Grade 11Algebra

find the sqare root of 4ab-2(a2-b2)i

Yash Shrivastava
11 Years agoGrade 11

1 Answer

Harsh Patodia
11 Years ago
Hi Student

Real Part of Above Complex Number is x= 4ab
Imaginary Part of Above Complex Number is y=-2(a2-b2)
Such that z= x+ iy
Now Calculating Square root of Complex Number
\sqrt{z} =\sqrt{x+iy} = ± ( \sqrt{(|z| + Re(z))/2 } - i\sqrt{(|z|-Re(z))/2}
where Re(z) means real part of z and |z| means modulus of z

Substitue the values in the above formula and you can get the answer.